(x-5000) * 126/100 = 6x/5
Answer by CubeyThePenguin(3113)

(x-5000) * 1.26 = 6x/5

6.3(x-5000) = 6x

6.3x - 31500 = 6x

0.3x = 31500

x = 105,000